OpenStudy (anonymous):

find all points on r=2sinθ where the tangent line is parallel to the ray θ=π/4

OpenStudy (anonymous):

(2sin pi/8,pi/8),(2 sin 5pi/8,5pi/8) are the answers if it helps!

OpenStudy (anonymous):

\[ \frac{dy}{dx}=\frac{\cos (\theta ) r(\theta )+\sin (\theta ) \frac{\partial r(\theta )}{\partial \theta }}{\cos (\theta ) \frac{\partial r(\theta )}{\partial \theta }-\sin (\theta ) r(\theta )}=\frac{4 \sin (\theta ) \cos (\theta )}{2 \cos ^2(\theta )-2 \sin ^2(\theta )} \] Set \[ \frac{4 \sin (\theta ) \cos (\theta )}{2 \cos ^2(\theta )-2 \sin ^2(\theta )} \]=1 and solve for \(\theta\) you find \[ \theta =\frac \pi 8\\ \theta =\frac {5\pi}8 \]

OpenStudy (anonymous):

\[ \frac{4 \sin (\theta ) \cos (\theta )}{2 \cos ^2(\theta )-2 \sin ^2(\theta )}=1\\ 4 \sin (\theta ) \cos (\theta )=2 \cos ^2(\theta )-2 \sin ^2(\theta )=2\cos(2 \theta)\\ 2 \sin(2\theta)= 2\cos(2 \theta)\\ \sin(2\theta)= \cos(2 \theta)\\ \] It is easy from now on

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

To finish your problem above, us \[ \sin(x)=\sin(y)\\ x=\frac \pi 4\\ x= \frac {5\pi}4\\ 0\le x \le 2\pi \]

ganeshie8 (ganeshie8):

Ahh nice :) I almost forgot about that dy/dx formula lol.. I took the hard way by converting it to cartesian and got \(\large y = 1 \pm\frac{1}{\sqrt{2}}\) ... again i need to switch it back to polar which is bit of pain :/ thanks @eliassaab for showing the work :)

OpenStudy (anonymous):

Thank you guys :)

OpenStudy (anonymous):

Why do you set the tangent line equal to one?

ganeshie8 (ganeshie8):

do u mean, why the "slope of tangent line" was set equal to 1 ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

read the question once again : ... the tangent line is parallel to the ray θ=π/4

ganeshie8 (ganeshie8):

whats the slope of ray : θ=π/4 ?

OpenStudy (anonymous):

oh alrights thanks!

ganeshie8 (ganeshie8):

:)